OpenStudy (anonymous):

if a and b are the roots of x^2+x+1=0, then find the equation with roots a^19 and b^7. the ans is x^2+x+=0...now try to find a way to get this ans

OpenStudy (anonymous):

\[a+b=-1\] \[ab=1\] solve a and b \[\huge \color{purple}{ \text{ welcome ROSE MATTHEW}}\]

OpenStudy (anonymous):

this is because in a quadratic equation with roots\[\alpha,\beta \text{ and form } ax^2+bx+c,\implies \alpha+\beta=\frac{ -b }{ a },\alpha \beta=\frac{ c }{ a }\] then we see that \[a(-1-a)=1\] \[a^2+a+1=0\]

OpenStudy (anonymous):

Let's first factor the given equation:\[\bf x^2+x+1=0 \implies x=\frac{ -1 \pm \sqrt{1-4(1)(1)} }{ 2 }=\frac{ -1 \pm \sqrt{3}i}{ 2 }\] So those are the two roots, now just make the polynomial as:\[\bf f(x)=(x-a^{19})(x-b^7)\]@rose.mathew

OpenStudy (anonymous):

@rose.mathew

OpenStudy (anonymous):

Show that \[ a^{19}=a\\ b^7=b \] and you are with your original equation.

OpenStudy (anonymous):

Thinking with groups, the set {1,a,b} is a group. Notice that \[ a=a\\ a^2=b\\ a^3=1\\ b=b\\ b^2=a\\ b^3=1\\ a b=1 \]